Grandmother has four grandchildren, two boys and two girls. When the ages of
the grandchildren are multiplied together, the product of their ages is 67 184.
Only one of the grandchildren is a teenager and the oldest grandchild is under
40. The difference in ages between the oldest and youngest grandchild is thirty
years.
Determine the ages of Grandmother's grandchildren.
What I have figured out so far is that the oldest grandchild will have an age between 31-39 and the youngest will be between 1-9 (since none of the ages can be zero-->this gives 0 as a product of all four ages). I believe that one of the other two remaining childrens' age is between 13-19. Is this information correct? How do I solve this word problem completely?

Answer by ankor@dixie-net.com(22740)

Grandmother has four grandchildren, two boys and two girls.
When the ages of the grandchildren are multiplied together, the product of their ages is 67 184.
Only one of the grandchildren is a teenager and the oldest grandchild is under
40.
The difference in ages between the oldest and youngest grandchild is thirty
years.
Determine the ages of Grandmother's grandchildren.
:
Also for the youngest, a limited number of single digits will have an integer quotient.
Assume the youngest is 8, and the oldest is 38
67184/8 = 8398; 8398/38 = 221, not enough for one teenager and one over 20
:
Try again, 6 doesn't give an integer, try 4
67184/4 = 16796; 16796/34 = 494; looks promising
Divide by some teenage numbers: 494/19 = 26, Bingo!
:
Grand-kid ages: 4, 19, 26, 34; check this 4*19*26*34 = 67184